# A note on division rings satisfying generalized rational identities with anti-automorphisms

### v2

@@
Abstract.
Let $D$ be a division ring with infinite center $F$; $\sigma$ be an anti-automorphism of $D$ and $m$ be a positive integer such that $\sigma^{m}\neq\mathrm{Id}$. In this paper, we show that if $D$ satisfies a $\sigma^{m}$-GRI, then $D$ is centrally finite.
Key words and phrases:
anti-automorphism, division ring, rational identity.
2020 Mathematics Subject Classification. 16K20, 16R50
1Faculty of Mathematics and Computer Science, University of Science,
Ho Chi Minh City, Vietnam
2Vietnam National University, Ho Chi Minh City, Vietnam
3Binh Duong University, Binh Duong Province, Vietnam
minhthuvohoang@gmail.com, vmtrang@bdu.edu.vn
1. Introduction and the main result
@@
A classic result states that if $D$ satisfies the polynomial identity, then $D$ is centrally finite (see [11]). To present this result and the extended results we aim to prove, we recall some concepts. Let $X=\{x_{1},x_{2},\cdots,x_{n}\}$ be a set of $n$ non-commutative indeterminates and $K$ be a field. We denote $K\langle X\rangle$ as the polynomial ring in indeterminates $X$ over $K$, and $R$ as a ring whose center contains $K$. A polynomial $f(X)\in K\langle X\rangle\backslash\{0\}$ is called a polynomial identity of $R$ if for every $n$-tuple $(a_{1},a_{2},\dots,a_{n})\in R^{n}$, we have $f(a_{1},a_{2},\dots,a_{n})=0$. In this case, we say $R$ satisfies the polynomial identity $f$, abbreviated as PI. If $R$ satisfies some polynomial identity, then we say $R$ satisfies a polynomial identity. Since we only work with division rings, we directly assume $R=D$ is a division ring and $K$ coincides with the center $F$ of $D$.
@@
In 1966, Amitsur proved that if the center $F$ is infinite and the division ring $D$ satisfies a generalized rational identity, then $D$ is centrally finite (see [2]). There are several extensions of this result. For example, in 1996, Chiba proved that if the center $F$ is infinite and the multiplicative group $D\{$$^{*}=D\backslash\{0\}$ contains a non-central proper subnormal subgroup satisfying a generalized rational identity, then $D$ is centrally finite (see [6]). Another direction is to consider identities corresponding to an anti-automorphism. Recall that a bijection $\sigma:D\to D,x\mapsto x^{\sigma},$ is called an anti-automorphism of $D$ if $(a+b)^{\sigma}=a^{\sigma}+b^{\sigma}$ and $(ab)^{\sigma}=b^{\sigma}a^{\sigma}$ for all $a,b\in D$. In the case where the order of $\sigma$ is $2$, $\sigma$ is called an involution of $D$. Let $m$ be a positive integer such that $\sigma^{i}\neq\mathrm{Id}$ for all $1\leq i\leq m$. For each $1\leq i\leq m$, denote $X^{\sigma^{i}}=\{x_{1}^{\sigma^{i}},x_{2}^{\sigma^{i}},\dots\}$ as the set of indeterminates indexed corresponding to each $\sigma^{i}$ and $X_{m}=\cup_{i=0}^{m}X^{\sigma^{i}}$. We denote $D(X_{m})$ as the division ring as mentioned above. Let $f(X_{m})$ be a nonzero element in $D(X_{m})$. If for every $n$-tuple $(r_{1},r_{2},\cdots,r_{n})\in D^{n}$ such that when
@@
then we say $f(X_{m})$ is a generalized rational identity of $D$ with respect to $\sigma^{m}$. In this case, we simply say that $D$ satisfies a $\sigma^{m}$-GRI $f$ or $f$ is a $\sigma^{m}$-GRI of $D$. The ”best” result is related to when $\sigma$ is an involution. In this case, $m=1$ and $\sigma$ will be denoted as $\star$. In more detail, it is proved that if the center $F$ is infinite and $D$ satisfies a $\star$-GRI, then $D$ is centrally finite (see [9]). The case of generalized polynomial identities with respect $\sigma$ receives even more attention (e.g., see the book [3] for a survey and the paper [18] for some additional interesting results on this topic). The objective of this paper is to prove the following result.
@@
2. Preliminaries
In short and simple terms, we prove the main result of this paper by induction on $n$, the number of indeterminates in the identity. In this section, we prove the case for $n=1$. To do it, we need to review some basic concepts and establish notation. Many results in this section may have proofs similar to those found in some reference documents, but none of the results are directly quoted from them. Therefore, for the convenience of the readers, we will provide proofs for the results if necessary. The main result we obtained in this section is Lemma  2.4, which states that if a division ring satisfies a blended linear identity in one variable, then this division ring is centrally finite.
We emphasize some notations that will be used throughout the paper as follows. Let $D$ be a division ring with center $F$, and $\sigma:D\to D,x\mapsto x^{\sigma},$ be an anti-automorphism over $F$, that is, $(a+b)^{\sigma}=a^{\sigma}+b^{\sigma},(ab)^{\sigma}=b^{\sigma}a^{\sigma}$ and $\alpha^{\sigma}=\alpha$ for every $a,b\in D$ and $\alpha\in F$. Consider $X=\{x_{1},x_{2},\dots\}$ as a countable set of non-commutative indeterminates. Suppose $m$ is an integer such that $\sigma^{i}\neq Id$ for all $1\leq i\leq m$. For each $1\leq i\leq m$, let $X^{\sigma^{i}}=\{x_{1}^{\sigma^{i}},x_{2}^{\sigma^{i}},\dots\}$ and $X_{m}=\bigcup_{i=0}^{m}X^{\sigma^{i}}$. The notation $F\langle X_{m}\rangle$ denotes the free algebra generated by $X_{m}$ over $F$, and $D\langle X_{m}\rangle=D*_{F}F\langle X_{m}\rangle$ is the free product of $D$ and $F\langle X_{m}\rangle$ over $F$. Each element of $D\langle X_{m}\rangle$ has the form
@@ 2. Preliminaries
For each monomial $P\in D\langle X_{m}\rangle$, we define the $\sigma^{m}$-degree of $P$ with respect to the indeterminate $x_{j}$, denoted $\sigma^{m}$-$\deg_{x_{j}}(P)$, as the number of occurrences of $x_{j}^{\sigma^{i}}$ in $P$ for some $0\leq i\leq m$. Then, the degree of $P$, denoted $\deg(P)$, is defined as
@@
We also define the $\sigma^{m}$-height of $P$ with respect to the indeterminate $x_{j}$, denoted $\sigma^{m}\text{-}ht_{x_{j}}(P)$, as the maximum value between $(\sigma^{m}\text{-}\deg_{x_{j}}(P)-1)$ and 0. Similarly, the height of $P$, denoted $ht(P)$, is defined as
@@
Suppose $f=f(x_{1},x_{2},\dots,x_{n},x_{1}^{\sigma},x_{2}^{\sigma},\dots,x_{n}^{\sigma},\dots,x_{1}^{\sigma^{m}},x_{2}^{\sigma^{m}},\dots,x_{n}^{\sigma^{m}})$ is a $\sigma^{m}$-GPI of $D$ with $\deg(f)=p$. Let $q$ be the number of indeterminates in $f$ that are not blended. Suppose $q>0$. Without loss of generality, we assume $f$ is not blended with respect to indeterminates $x_{1},x_{2},\cdots,x_{q}$. Set
@@
$f_{2j-1}$ $=f_{2j-2}(x_{1},x_{2},\dots,x_{j-1},0,$
$\quad x_{j+1},\cdots,x_{n},\cdots,x_{1}^{\sigma^{m}},x_{2}^{\sigma^{m}},\dots,x_{j-1}^{\sigma^{m}},0,x_{j+1}^{\sigma^{m}},\cdots,x_{n}^{\sigma^{m}}),$
@@
We continue to consider $f=\sum_{i=1}^{\ell}P_{i}\in D\langle X_{m}\rangle$ where $P_{1},P_{2},\dots,P_{\ell}$ are monomials. Then, $f$ is called $\sigma^{m}$-multilinear (or shortly, $\sigma^{m}$-linear) of degree $n$ on indeterminates $x_{1},x_{2},\cdots,x_{n}$ if each monomial $P_{i}$ in $f$ satisfies the following properties:
  1. (1)
@@
  2. (2)
    $\sigma^{m}\text{-}\deg_{x_{j}}(P_{i})=1$ for all $x_{j}\in\{x_{1},x_{2},\cdots,x_{n}\}$.
@@
Given $f=\sum_{i=1}^{\ell}P_{i}\in D\langle X_{m}\rangle$ is a blended polynomial in indeterminates $x_{1},x_{2},\dots,x_{n}$ and has height $t$. Let $P$ be a monomial in $f$ such that $ht(P)=t$. Then, $\Delta_{1}^{(p_{1}-1)}\dots\Delta_{n}^{(p_{n}-1)}f$ is a $\sigma^{m}$-linear polynomial in $\sum_{j=1}^{n}p_{j}$ indeterminates, where $p_{j}=\sigma^{m}\text{-}\deg_{x_{j}}(P)$ $(1\leq j\leq n)$.
Proof.  If $f$ has height $t=0$, then $p_{j}=1$ and $\Delta_{j}f=f$ is $\sigma^{m}$-linear for every $1\leq j\leq n$, thus the lemma is proved. Consider the case $t>0$ and $P$ is a monomial in $f$ such that $\sigma^{m}\text{-ht}(P)=t$. With $p_{j}=\sigma^{m}\text{-}\deg_{x_{j}}(P)$, we have
$\Delta_{1}^{(p_{1}-1)}\dots\Delta_{n}^{(p_{n}-1)}f=\sum\Delta_{1}^{(p_{1}-1)}\dots\Delta_{n}^{(p_{n}-1)}P_{i}.$
Choose $j_{0}$ such that $\sigma^{m}\text{-}\deg_{x_{j_{0}}}(P)=p_{j_{0}}>1$ and $\deg_{x_{i}}(P)=p_{i}=1$ in which $1\leq j_{0}<i\leq n$. Then
$\Delta_{1}^{(p_{1}-1)}\dots\Delta_{n}^{(p_{n}-1)}P=\Delta_{1}^{(p_{1}-1)}\dots\Delta_{j_{0}}^{(p_{j_{0}}-1)}P.$
@@
Each polynomial $h_{u}$ is a sum of $C_{p_{j_{0}}}^{u}$ monomials $P_{u_{k}}$, where each monomial $P_{u_{k}}$ maps to $P$ under the projection $x_{n+1}\to x_{j_{0}}$ and $\sigma^{m}\text{-}\deg_{x_{j_{0}}}(P_{u_{k}})=p_{j_{0}}-u$. Hence,
@@
monomials, each of which maps to $P$ under the projection $x_{e}\mapsto x_{j_{0}}$ (and here $e=n+1,n+2\cdots,n+(p_{j_{0}}-1$). Furthermore, we have
@@
Let $P_{i_{0}}\neq P$ be a monomial of $f$. If there exists $1\leq j_{k}\leq n$ such that $\sigma$-$\deg_{x_{{j_{k}}}}(P_{i_{0}})<p_{j_{k}}=\sigma$-$\deg_{x_{j_{k}}}(P)$, then $\Delta_{{j_{k}}}^{(p_{j_{k}}-1)}P_{i_{0}}=0$. So
@@
Proof.  By Corollary 2.3, we assume that $f$ is blended $\sigma^{m}$-linear. By the hypothesis, $f$ can be represented as
@@
Suppose conversely, for all $r\in D$, $\sum_{i=0}^{m}g_{i}(r^{\sigma^{i}})=0.$ Consider the polynomial $g_{0}(x)=\sum^{n_{0}}_{j=1}a_{0j}xb_{0j}$ in two cases $n_{0}=0$ and $n_{0}>0$.In the case. First, we consider the case $n_{0}>0$. If the set $\{a_{01},a_{02},\cdots,a_{0n_{0}}\}$ of $n_{0}$ elements is linearly independent over $F$, then the set of $2n_{0}$ polynomials
$a_{01},a_{02},a_{03},\cdots,a_{0n_{0}},a_{mn_{n}}xa_{01},a_{mn_{m}}xa_{02},a_{mn_{m}}xa_{03},\cdots,a_{mn_{m}}xa_{0n_{0}}$
@@
where $\alpha_{j},\beta_{j}\in F$ and $j\in\{1;2;\cdots;n_{0}\}$. Then, we have
@@
Since $\{a_{01},a_{02},\cdots,a_{0n_{0}}\}$ is linearly independent over $F$, we have the following: $\alpha_{1}=\alpha_{2}=\cdots=\ \alpha_{n}=0.$ Furthermore, as $\{a_{01},a_{02},\cdots,a_{0n_{0}}\}$ is linearly independent over $F$, the set
$\{a_{mn_{n}}\otimes_{F}a_{01};a_{mn_{m}}\otimes_{F}a_{02};\cdots;a_{mn_{m}}\otimes_{F}a_{0n_{0}}\}$
@@
are also linearly independent over $F$. According to [3, Remark 6.1.1], there exists an isomorphism $\phi:D\otimes_{F}D\to DxD$ such that $a\otimes_{F}\mapsto axb$. Since
@@
$f=f(x^{\sigma},\cdots,x^{\sigma^{m}})$ $=\sum_{i=0}^{m}g_{i}(x^{\sigma^{i}})$
@@
$0$ $=f(r^{\sigma},\cdots,r^{\sigma^{m}})$
@@
3. Proof of the main result

### v3

@@
Abstract
Let $D$ be a division ring with infinite center $F$, $\sigma$ be an anti-automorphism of $D$ and $m$ be a positive integer such that $\sigma^{m}\neq\mathrm{Id}$. In this paper, we show that if $D$ satisfies a $\sigma^{m}$-GRI, then $D$ is centrally finite.
keywords:
anti-automorphism, division ring, rational identity
[…]
1 Introduction and the main result
@@
A classical result states that if $D$ satisfies a polynomial identity, then $D$ is centrally finite (see [11]). To present this result and the extended results we aim to prove, we recall some concepts. Let $X=\{x_{1},x_{2},\dots,x_{n}\}$ be a set of $n$ non-commutative indeterminates and $K$ be a field. We denote $K\langle X\rangle$ as the polynomial ring in indeterminates $X$ over $K$, and $R$ as a ring whose center contains $K$. A polynomial $f(X)\in K\langle X\rangle\backslash\{0\}$ is called a polynomial identity of $R$ if for every $n$-tuple $(a_{1},a_{2},\dots,a_{n})\in R^{n}$, we have $f(a_{1},a_{2},\dots,a_{n})=0$. In this case, we say $R$ satisfies the polynomial identity $f$, abbreviated as PI. If $R$ satisfies some polynomial identity, then we say $R$ satisfies a polynomial identity. Since we only work with division rings, we directly assume $R=D$ is a division ring and $K$ coincides with the center $F$ of $D$.
@@
In 1966, Amitsur proved that if the center $F$ is infinite and the division ring $D$ satisfies a generalized rational identity, then $D$ is centrally finite (see [2]). There are several extensions of this result. For example, in 1996, Chiba proved that if the center $F$ is infinite and the multiplicative group $D\{$$^{*}=D\backslash\{0\}$ contains a non-central proper subnormal subgroup satisfying a generalized rational identity, then $D$ is centrally finite (see [6]). Another direction is to consider identities corresponding to an anti-automorphism. Recall that a bijection $\sigma:D\to D,x\mapsto x^{\sigma},$ is called an anti-automorphism of $D$ if $(a+b)^{\sigma}=a^{\sigma}+b^{\sigma}$ and $(ab)^{\sigma}=b^{\sigma}a^{\sigma}$ for all $a,b\in D$. In the case where the order of $\sigma$ is $2$, $\sigma$ is called an involution of $D$. Let $m$ be a positive integer such that $\sigma^{i}\neq\mathrm{Id}$ for all $1\leq i\leq m$. For each $1\leq i\leq m$, denote $X^{\sigma^{i}}=\{x_{1}^{\sigma^{i}},x_{2}^{\sigma^{i}},\dots\}$ as the set of indeterminates indexed corresponding to each $\sigma^{i}$ and $X_{m}=\cup_{i=0}^{m}X^{\sigma^{i}}$. We denote $D(X_{m})$ as the division ring as mentioned above. Let $f(X_{m})$ be a nonzero element in $D(X_{m})$. If for every $n$-tuple $(r_{1},r_{2},\dots,r_{n})\in D^{n}$ such that when
@@
then we say $f(X_{m})$ is a generalized rational identity of $D$ with respect to $\sigma^{m}$. In this case, we simply say that $D$ satisfies a $\sigma^{m}$-GRI $f$ or $f$ is a $\sigma^{m}$-GRI of $D$. The “best” result is related to when $\sigma$ is an involution. In this case, $m=1$ and $\sigma$ will be denoted as $\star$. In more detail, it is proved that if the center $F$ is infinite and $D$ satisfies a $\star$-GRI, then $D$ is centrally finite (see [9]). The case of generalized polynomial identities with respect to $\sigma$ receives even more attention (e.g., see the book [3] for a survey and the paper [18] for some additional interesting results on this topic). The objective of this paper is to prove the following result.
@@
2 Preliminaries
In short and simple terms, we prove the main result of this paper by induction on $n$, the number of indeterminates in the identity. In this section, we prove the case for $n=1$. To do it, we need to review some basic concepts and establish notation. Many results in this section may have proofs similar to those found in some reference documents, but none of the results are directly quoted from them. Therefore, for the convenience of the readers, we will provide proofs for the results if necessary. The main result we obtained in this section is Lemma  2.5, which states that if a division ring satisfies a blended linear identity in one variable, then this division ring is centrally finite.
We emphasize some notations that will be used throughout the paper as follows. Let $D$ be a division ring with center $F$, and $\sigma:D\to D,x\mapsto x^{\sigma},$ be an anti-automorphism over $F$, that is, $(a+b)^{\sigma}=a^{\sigma}+b^{\sigma},(ab)^{\sigma}=b^{\sigma}a^{\sigma}$ and $\alpha^{\sigma}=\alpha$, for every $a,b\in D$ and $\alpha\in F$. Consider $X=\{x_{1},x_{2},\dots\}$ as a countable set of non-commutative indeterminates. Suppose $m$ is an integer such that $\sigma^{i}\neq Id$ for all $1\leq i\leq m$. For each $1\leq i\leq m$, let $X^{\sigma^{i}}=\{x_{1}^{\sigma^{i}},x_{2}^{\sigma^{i}},\dots\}$ and $X_{m}=\bigcup_{i=0}^{m}X^{\sigma^{i}}$. The notation $F\langle X_{m}\rangle$ denotes the free algebra generated by $X_{m}$ over $F$, and $D\langle X_{m}\rangle=D*_{F}F\langle X_{m}\rangle$ is the free product of $D$ and $F\langle X_{m}\rangle$ over $F$. Each element of $D\langle X_{m}\rangle$ has the form
@@ 2 Preliminaries
For each monomial $P\in D\langle X_{m}\rangle$, we define the $\sigma^{m}$-degree of $P$ with respect to the indeterminate $x_{j}$, denoted $\sigma^{m}$-$\deg_{x_{j}}(P)$, as the number of occurrences of $x_{j}^{\sigma^{i}}$ in $P$ for some $0\leq i\leq m$. Then, the degree of $P$, denoted by $\deg(P)$, is defined as
@@
We also define the $\sigma^{m}$-height of $P$ with respect to the indeterminate $x_{j}$, denoted $\sigma^{m}\text{-}ht_{x_{j}}(P)$, as the maximum value between $(\sigma^{m}\text{-}\deg_{x_{j}}(P)-1)$ and 0. Similarly, the height of $P$, denoted by $ht(P)$, is defined as
@@
Suppose
$f=f(x_{1},x_{2},\dots,x_{n},x_{1}^{\sigma},x_{2}^{\sigma},\dots,x_{n}^{\sigma},\dots,x_{1}^{\sigma^{m}},x_{2}^{\sigma^{m}},\dots,x_{n}^{\sigma^{m}})$
is a $\sigma^{m}$-GPI of $D$ with $\deg(f)=p$. Let $q$ be the number of indeterminates in $f$ that are not blended. Suppose $q>0$. Without loss of generality, we assume $f$ is not blended with respect to indeterminates $x_{1},x_{2},\cdots,x_{q}$. Set
@@
$f_{2j-1}=f_{2j-2}(x_{1},x_{2},\dots,x_{j-1},0,x_{j+1},\cdots,x_{n},\cdots,x_{1}^{\sigma^{m}},x_{2}^{\sigma^{m}},\dots,x_{j-1}^{\sigma^{m}},0,x_{j+1}^{\sigma^{m}},\cdots,x_{n}^{\sigma^{m}}),$
@@
We continue to consider $f=\sum_{i=1}^{\ell}P_{i}\in D\langle X_{m}\rangle$ where $P_{1},P_{2},\dots,P_{\ell}$ are monomials. Then, $f$ is called $\sigma^{m}$-multilinear (or shortly, $\sigma^{m}$-linear) of degree $n$ on indeterminates $x_{1},x_{2},\dots,x_{n}$ if each monomial $P_{i}$ in $f$ satisfies the following properties:
  1. 1.
@@
  2. 2.
    $\sigma^{m}\text{-}\deg_{x_{j}}(P_{i})=1$ for all $x_{j}\in\{x_{1},x_{2},\dots,x_{n}\}$.
@@
Given $f=\sum_{i=1}^{\ell}P_{i}\in D\langle X_{m}\rangle$ is a blended polynomial in indeterminates $x_{1},x_{2},\dots,x_{n}$ and has height $t$. Let $P$ be a monomial in $f$ such that $ht(P)=t$. Then, we have that $\Delta_{1}^{(p_{1}-1)}\cdots\Delta_{n}^{(p_{n}-1)}f$ is a $\sigma^{m}$-linear polynomial in $\sum_{j=1}^{n}p_{j}$ indeterminates, where $p_{j}=\sigma^{m}\text{-}\deg_{x_{j}}(P)$ $(1\leq j\leq n)$.
Proof 2.3.
If $f$ has height $t=0$, then $p_{j}=1$ and $\Delta_{j}f=f$ is $\sigma^{m}$-linear for every $1\leq j\leq n$, thus the lemma is proved. Consider the case $t>0$ and $P$ is a monomial in $f$ such that $\sigma^{m}\text{-ht}(P)=t$. With $p_{j}=\sigma^{m}\text{-}\deg_{x_{j}}(P)$, we have
$\Delta_{1}^{(p_{1}-1)}\cdots\Delta_{n}^{(p_{n}-1)}f=\sum\Delta_{1}^{(p_{1}-1)}\cdots\Delta_{n}^{(p_{n}-1)}P_{i}.$
Choose $j_{0}$ such that $\sigma^{m}\text{-}\deg_{x_{j_{0}}}(P)=p_{j_{0}}>1$, $\deg_{x_{i}}(P)=p_{i}=1$ and $1\leq j_{0}<i\leq n$. Then
$\Delta_{1}^{(p_{1}-1)}\cdots\Delta_{n}^{(p_{n}-1)}P=\Delta_{1}^{(p_{1}-1)}\cdots\Delta_{j_{0}}^{(p_{j_{0}}-1)}P.$
@@
Each polynomial $h_{u}$ is a sum of $C_{p_{j_{0}}}^{u}$ monomials $P_{u_{k}}$, where each monomial $P_{u_{k}}$ maps to $P$ under the projection $x_{n+1}\mapsto x_{j_{0}}$ and $\sigma^{m}\text{-}\deg_{x_{j_{0}}}(P_{u_{k}})=p_{j_{0}}-u$. Hence,
@@
monomials, each of which maps to $P$ under the projection $x_{e}\mapsto x_{j_{0}}$ (and here we denote $e=n+1,n+2,\dots,n+(p_{j_{0}}-1$)). Furthermore, we have
@@
Let $P_{i_{0}}\neq P$ be a monomial of $f$. If there exists $1\leq j_{k}\leq n$ satisfying the inequality $\sigma$-$\deg_{x_{{j_{k}}}}(P_{i_{0}})<p_{j_{k}}=\sigma$-$\deg_{x_{j_{k}}}(P)$, then $\Delta_{{j_{k}}}^{(p_{j_{k}}-1)}P_{i_{0}}=0$. So
@@
Proof 2.6.
By Corollary 2.4, we assume that $f$ is blended $\sigma^{m}$-linear. By the hypothesis, $f$ can be represented as
@@
Suppose conversely that we have $\sum_{i=0}^{m}g_{i}(r^{\sigma^{i}})=0,$ for all $r\in D$. Consider the polynomial $g_{0}(x)=\sum^{n_{0}}_{j=1}a_{0j}xb_{0j}$ in two cases $n_{0}=0$ and $n_{0}>0$. First, we consider the case $n_{0}>0$. If the set $\{a_{01},a_{02},\dots,a_{0n_{0}}\}$ of $n_{0}$ elements is linearly independent over $F$, then the set of $2n_{0}$ polynomials
$a_{01},a_{02},a_{03},\dots,a_{0n_{0}},a_{mn_{n}}xa_{01},a_{mn_{m}}xa_{02},a_{mn_{m}}xa_{03},\dots,a_{mn_{m}}xa_{0n_{0}}$
@@
where $\alpha_{j},\beta_{j}\in F$ and $j\in\{1,2,\dots,n_{0}\}$. Then, we have
@@
Since $\{a_{01},a_{02},\dots,a_{0n_{0}}\}$ is linearly independent over $F$, we have the following relation: $\alpha_{1}=\alpha_{2}=\cdots=\ \alpha_{n}=0.$ Furthermore, as $\{a_{01},a_{02},\dots,a_{0n_{0}}\}$ is linearly independent over $F$, the set
$\{a_{mn_{n}}\otimes_{F}a_{01},a_{mn_{m}}\otimes_{F}a_{02},\dots,a_{mn_{m}}\otimes_{F}a_{0n_{0}}\}$
@@
are also linearly independent over $F$. According to [3, Remark 6.1.1], there exists an isomorphism $\phi:D\otimes_{F}D\to DxD$ such that $a\otimes_{F}b\mapsto axb$. Since
@@
$f=f(x^{\sigma},\dots,x^{\sigma^{m}})$ $=\sum_{i=0}^{m}g_{i}(x^{\sigma^{i}})$
@@
$0$ $=f(r^{\sigma},\dots,r^{\sigma^{m}})$
@@
3 Proof of the main result